OpenStudy (anonymous):

A triangle formed by the x-axis, y-axis and the line y = ax + b, if a = -b then what is the volume of the cone generated by rotating the triangle around y-axis?

OpenStudy (anonymous):

the answer is pi/3 a @Hezum

OpenStudy (anonymous):

Ok thanks rose, but how? can you tell me the solution?

OpenStudy (anonymous):

Or if you can attach here the screenshot.

OpenStudy (anonymous):

the height is b as him said - the intercept on the x axis is -b1/2 so the volume is (1/3) pi r^2 h = (1/3) pi 1^2 * -b =(1/3)* pi * a = (pi/3) * a which is b) on the list

OpenStudy (anonymous):

from where u get this question?

OpenStudy (anonymous):

I got it from Iba Cs past entry test

OpenStudy (anonymous):

ok same question i posted

OpenStudy (anonymous):

u r going to gave ibas test?

OpenStudy (anonymous):

yes im going for it.

OpenStudy (anonymous):

you too?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

where are u from?

OpenStudy (anonymous):

and how u are preparing its on coming sunday

OpenStudy (anonymous):

Im from Karachi Malir cantt

OpenStudy (anonymous):

Well honestly speaking im not preapring well. just 5 days left and i haven't done anything special for it.

OpenStudy (anonymous):

Now i have decided to solve the past papers that are uploaded on the website

OpenStudy (anonymous):

I tooo :D

OpenStudy (anonymous):

you from where?

OpenStudy (anonymous):

Oh i just realized i dint thank you for it. So, thank youu so much rose. This question was cooking me so badly

OpenStudy (anonymous):

if u want to take any help from me so ur welcome

OpenStudy (anonymous):

Okay sstay connected.

OpenStudy (anonymous):

you on fb? no im just asking because im not using this web from my personal account and it is laging like anything,

OpenStudy (anonymous):

hmm u too

OpenStudy (anonymous):

Well, i will deal with the essay part. no preaparation for it. im focused on Mathematics.

OpenStudy (anonymous):

same here:D

OpenStudy (anonymous):

No gmail here. :p

OpenStudy (anonymous):

Best of luck for sunday then.

OpenStudy (anonymous):

You don't have to prepare for Essay, as there is no essay part in BS test

OpenStudy (anonymous):

Woah! thats good then. :)

OpenStudy (anonymous):

What's your fb Hezum?

OpenStudy (anonymous):

its hassaan.ahmed@outlook.com

OpenStudy (anonymous):

guys no personal information on ques...pm each other for such things

OpenStudy (anonymous):

ok then give the one you have on outlook

OpenStudy (anonymous):

@rosebird

OpenStudy (anonymous):

siraj its hassaan.ahmed95@outlook.com

OpenStudy (anonymous):

Added you @Hezum , please stay connected :P P.s: @rosebird what's your gmail?

OpenStudy (anonymous):

okay boy :p

OpenStudy (anonymous):

LOL :D

OpenStudy (anonymous):

hmm best of luck guyz for test

OpenStudy (anonymous):

Same to yeww! :)

OpenStudy (anonymous):

Best of luck toy you all. hope we'll meet there with happy faces after the test duration :p

OpenStudy (anonymous):

inshallah :)

OpenStudy (anonymous):

Aameen ^_^

OpenStudy (anonymous):

Inshallah :)

OpenStudy (anonymous):

pray that we clear the test n interview aswell

OpenStudy (anonymous):

I never expected that i'd meet IBA candidates at openstudy, LOL :D

OpenStudy (anonymous):

I too never expected. Now to study with you guyz there in IBA.

OpenStudy (anonymous):

now I wish*

OpenStudy (anonymous):

yeh same here i posted this question before u guyz then i saw after me u boyh also posted it so i ask hezum from where u get this question in this way i meet u guyz

OpenStudy (anonymous):

I have some sources through which you guys can prepare for the test, i'll send em to you both.

OpenStudy (anonymous):

Thats so kind from your side siraj. Thanks boy!

OpenStudy (anonymous):

thanks @sirajawahid

OpenStudy (anonymous):

No need to mention @Hezum and @rosebird :)

OpenStudy (anonymous):

:)